i need to determine if this equation is true or false and if its false i need to correct the first part If |y+8|+2=6, then y+8=+_4 the plus sign is showed to be over the underline but i can't get my computer to line it up
Answer by ewatrrr(24785)

If |y+8|+2=6, ⇒ |y+8|= 4
and Yes, y+8 = 4 0r y+8 = -4
y + 8 = ± 4